Two RSM students are standing 20ft apart and spot Mrs Rifkin. If they immediately run opposite direction with speeds of 300 ft/min and 200 ft/min, how many hours will pass until they are 3,000 ft apart?
Answer by ankor@dixie-net.com(22740)

Two RSM students are standing 20ft apart and spot Mrs Rifkin.
If they immediately run opposite direction with speeds of 300 ft/min and 200 ft/min, how many hours will pass until they are 3,000 ft apart?
:
Let t = no. of minutes for them to be 3000 ft apart
300t + 200t + 20 = 3000
500t = 3000 - 20
t = 2980/500
t = 5.96 min
Change to hrs
5.96/60 = .0933 hrs